# The Information Dynamics of Generative Diffusion

### v1

@@ 3 Information theory
Consider a game of Twenty Questions where an interrogator player may ask twenty binary questions concerning a set to an ”oracle” player in order to gradually reveal the identity of a predetermined element $\mathbf{y}$$^{*}$ in a finite set $\Omega=\{{\mathbf{y}}_{1},\dots,{\mathbf{y}}_{N_{0}}\}$ with $N_{0}$ elements. We denote the size of the possible set $\Omega_{j-1}(a_{1:j-1})$ after j1j-1 questions as $N_{j-1}(a_{1:j-1})$. The answer $a_{j}$ to the $j$-th question $q_{j}$ then divides the set into two possible subsets with sizes $N^{1}_{j}(a_{1:j-1})=N_{j}(a_{j}=1,a_{1:j-1})$ and $N^{0}_{j}(a_{1:j-1})=N_{j-1}(a_{1:j-1})-N_{j}(a_{j}=0,a_{1:j-1})$. Assuming a fixed set of questions, the expected uncertainty experienced by the player at after the $j$-th question can be quantified by the conditional entropy:
@@ 5.1 Branching paths of fixed-points and spontaneous symmetry breaking
We denote the set of fixed-points at time $t$ as $\Psi_{t}$. The solutions of this fixed-point equation can be organized in a set $\Omega$ of piecewise continuous paths $\bm{\gamma}:\mathds{R}^{+}\rightarrow\mathds{R}^{d}\in\Omega$. To remove ambiguities, we assume that, if $\bm{\gamma}(\tau)$ is discontinuous at $\tau_{0}$, then the one-sided limit exists and it is equal to $\lim_{t\rightarrow\tau^{+}}\bm{\gamma}(\tau)=\operatorname*{arg\,min}_{\bm{x}\in\Psi_{t}}\left[\left\lVert\bm{x}-\bm{\gamma}(\tau)\right\rVert\right]$. We know that $\lim_{t\rightarrow\infty}\bm{\gamma}(t)=\bm{0}$ for all paths since the zero vector is the only fixed-point of the score of the asymptotic Gaussian distribution. From this, it follows that any two paths $\bm{\gamma}_{1}(t)$ and $\bm{\gamma}_{2}(t)$ can be proven to overlap for a finite range of time, meaning that $\bm{\gamma}_{1}(t)=\bm{\gamma}_{2}(t)$ if $t\geq\tau_{12}\in\mathds{R}^{+}$. We refer to $\tau_{1,2}$ as the branching time of the two paths. The branching time of two paths of fixed points can roughly be interpreted as a decision time in the generative process, where the sample will be ’pushed’ by the noise in either one or the other path during the reverse dynamics. It is therefore insightful to study the behavior of the paths at the branching times. In general, this can happen if there is a discontinuous jump in a path $\gamma(t)$. Perhaps more interestingly, two paths can also branch continuously at a finite time. This can be studied by analyzing the Jacobian matrix of the score function:
@@
If we approximate the interference function with constant $\phi$ using a zero-th order Taylor expansion, Eq. 26 becomes the self-consistency equation of a Curie-Weiss model of magnetism, with temperature $T=\sigma^{2}(t)$ and external magnetic field $\phi$. The solutions of this equation can be visualized as intersection points between a straight line and a hyperbolic tangent (see (Ambrogioni, 2024b) and (Raya & Ambrogioni, 2023) for a detail analysis). When $\phi$ is finite, the system transitions discontinuously from one to two fixed-points, which corresponds to a first-order phase transition in the magnetic system. However, the size of the discontinuity vanishes when $\phi=0$, when there is an exact symmetry between the two data-points (see Fig. 1). This gives raise to a so called critical phase transition, where a single fixed-point at $x^{*}=0$ continuously splits into two paths $\bm{x}_{1}(t)$ and $x_{2}(t)$ with $\bm{x}_{1,2}(t-t_{c})\sim\pm(t-t_{c})^{1/2}$ for $t\rightarrow 0$. The loss of stability of the fixed-point at the origin corresponds to the vanishing of the quadratic well around the point:

### v2

@@ 3 Information theory
Consider a game of Twenty Questions where an interrogator player may ask twenty binary questions concerning a set to an ”oracle” player in order to gradually reveal the identity of a predetermined element $\mathbf{y}$$^{*}$ in a finite set $\Omega=\{{\mathbf{y}}_{1},\dots,{\mathbf{y}}_{N_{0}}\}$ with $N_{0}$ elements. We denote the size of the possible set $\Omega_{j-1}(a_{1:j-1})$ after j1j-1 questions as $N_{j-1}(a_{1:j-1})$. The answer $a_{j}$ to the $j$-th question $q_{j}$ then divides the set into two possible subsets with sizes $N^{1}_{j}(a_{1:j-1})=N_{j}(a_{j}=1,a_{1:j-1})$ and $N^{0}_{j}(a_{1:j-1})=N_{j-1}(a_{1:j-1})-N_{j}(a_{j}=0,a_{1:j-1})$. Assuming a fixed set of questions, the expected uncertainty experienced by the player after the $j$-th question can be quantified by the conditional entropy:
@@ 4 Generative information transfer in score matching diffusion
4.1 Conditional entropy production as optimal error
The conditional entropy rate quantifies the instantaneous generative information transfer at any given moment in time. It can be shown (see (Stancevic et al., 2025)) that this quantity is closely connected to the optimal denoising squared error, which is the variance of the denoising distribution:
$\dot{\mathbf{H}}[\bm{y}|\bm{x}_{t}]=\frac{1}{2}\frac{\nu^{2}(t)}{\sigma^{4}(t)}\mathbb{E}_{\bm{y},\bm{x}_{t}}\left[\text{var}(\bm{y}\mid\bm{x}_{t})\right]~.$ (17)
Intuitively, this means that the information rate is directly related to the denoising uncertainty at a given time.
Using this relation, we can now re-express the denoising score matching formula in Eq. 9 in terms of the conditional entropy rate:
$\mathbb{E}_{\bm{x}_{t}}\left[\left\lVert\mathbb{E}_{{\mathbf{z}}\mid\bm{x}_{t}}\left[{\mathbf{z}}\right]-s(\bm{x}_{t};\theta)\right\rVert^{2}\right]+\frac{2\sigma^{4}(t)}{\nu^{2}(t)}\dot{\mathbf{H}}[\bm{y}|\bm{x}_{t}]=\mathbb{E}_{{\mathbf{z}},{\mathbf{y}}}\left[\left\lVert{\mathbf{z}}-s(\bm{y}+\sigma(t){\mathbf{z}};\theta)\right\rVert^{2}\right]~,$ (18)
which implies that the entropy rate can be estimated from the training loss if we assume that the network is well trained.
@@ 5.1 Branching paths of fixed-points and spontaneous symmetry breaking
We denote the set of fixed-points at time $t$ as $\Psi_{t}$. The solutions of this fixed-point equation can be organized in a set $\Omega$ of piecewise continuous paths $\bm{\gamma}:\mathds{R}^{+}\rightarrow\mathds{R}^{d}\in\Omega$. To remove ambiguities, we assume that, if $\bm{\gamma}(\tau)$ is discontinuous at $\tau_{0}$, then the one-sided limit exists and it is equal to $\lim_{t\rightarrow\tau^{+}}\bm{\gamma}(\tau)=\operatorname*{arg\,min}_{\bm{x}\in\Psi_{t}}\left[\left\lVert\bm{x}-\bm{\gamma}(\tau)\right\rVert\right]$. We know that $\lim_{t\rightarrow\infty}\bm{\gamma}(t)=\bm{0}$ for all paths since the zero vector is the only fixed-point of the score of the asymptotic Gaussian distribution. Any two paths $\bm{\gamma}_{1}(t)$ and $\bm{\gamma}_{2}(t)$ can be proven to overlap for a finite range of time, meaning that $\bm{\gamma}_{1}(t)=\bm{\gamma}_{2}(t)$ if $t\geq\tau_{12}\in\mathds{R}^{+}$ (this follows from the results in (Carreira-Perpinan, 2002; Carreira-Perpinán & Williams, 2003; Améndola et al., 2020) on the number of modes of mixture of normal distributions). We refer to $\tau_{1,2}$ as the branching time of the two paths. The branching time of two paths of fixed points can roughly be interpreted as a decision time in the generative process, where the sample will be ’pushed’ by the noise in either one or the other path during the reverse dynamics. It is therefore insightful to study the behavior of the paths at the branching times. In general, this can happen if there is a discontinuous jump in a path $\gamma(t)$. Perhaps more interestingly, two paths can also branch continuously at a finite time. This can be studied by analyzing the Jacobian matrix of the score function:
@@
If we approximate the interference function with constant $\phi$ using a zero-th order Taylor expansion, Eq. 28 becomes the self-consistency equation of a Curie-Weiss model of magnetism, with temperature $T=\sigma^{2}(t)$ and external magnetic field $\phi$. The solutions of this equation can be visualized as intersection points between a straight line and a hyperbolic tangent (see (Ambrogioni, 2024b) and (Raya & Ambrogioni, 2023) for a detail analysis). When $\phi$ is finite, the system transitions discontinuously from one to two fixed-points, which corresponds to a first-order phase transition in the magnetic system. However, the size of the discontinuity vanishes when $\phi=0$, when there is an exact symmetry between the two data-points (see Fig. 1). This gives rise to a so called critical phase transition, where a single fixed-point at $x^{*}=0$ continuously splits into two paths $\bm{x}_{1}(t)$ and $x_{2}(t)$ with $\bm{x}_{1,2}(t-t_{c})\sim\pm(t-t_{c})^{1/2}$ for $t\rightarrow 0$. The loss of stability of the fixed-point at the origin corresponds to the vanishing of the quadratic well around the point: